Assume that for the past 10 years the growth rate in A Ltd's dividends per share has been 10% p.a. Dividends are paid once a year. Also assume that this growth rate is to be maintained indefinitely. The latest dividend of 90 cents was paid yesterday. The required rate of return is 16% p.a. What is the value of A's shares?Assume that for the past 10 years the growth rate in A Ltd's dividends per share has been 10% p.a. Dividends are paid once a year. Also assume that this growth rate is to be maintained indefinitely. The latest dividend of 90 cents was paid yesterday. The required rate of return is 16% p.a. What is the value of A's shares?
Answer #1

value of share = D0(1+g) / K - g

where D0 = last dividend

g = growth rate

k = required return

so value of A's share = 0.9(1+10%) / (16% - 10%)

(90 cents = $0.9)

= 0.99 / 6%

=$16.5